Solve the equation.x+3/6=1- x+6/7
1
Expert's answer
2020-08-30T17:59:09-0400

x+3/6=1- x+6/7

2*x+1/2=1+6/7

2*x=13/7-1/2

2*x=19/14

x=19/28